OpenStudy (anonymous):

PLz help....((x-7)/2)+((x+5)/2)

OpenStudy (pfenn1):

\[\frac{x-7}{2}+\frac{x+5}{2}=\frac{x-7+x+5}{2}=\frac{2x-2}{2}=x-1\]

OpenStudy (anonymous):

Thanks for showing how to do it.

OpenStudy (pfenn1):

you're welcome